A) Solve: Log( 0.000001) = 2x-5 B) Solve: x3/5 = -27
Answer #1

- 2x-5 solution - A) Log (0.000001) log (106) = 2x-5 Joglo 10 = 2x-5 [ 3-6=2x-5 Ľ : log10=1] 6 A = B AB log Blog A - 2x - al